OpenStudy (idealist10):

How to integrate sqrt(2y-y^2)*y dy?

OpenStudy (idealist10):

@ganeshie8 @Luigi0210 @paki @abb0t @Destinymasha

OpenStudy (idealist10):

\[\int\limits \sqrt{2y-y^2}y dy\]

OpenStudy (idealist10):

@hartnn

OpenStudy (perl):

I would start by completing the square under the radical

OpenStudy (perl):

\[\int\limits_{}^{} y*\sqrt{2y-y^2}dy= \int\limits_{}^{} y*\sqrt{1-(y-1)^2}dy\]

OpenStudy (perl):

then do a u-substitution, u = y-1

OpenStudy (idealist10):

How did you get 1-(y-1)^2?

OpenStudy (perl):

2y - y^2 = -y^2 + 2y = - ( y^2 - 2y ) complete the square = - ( y^2 - 2y + 1 - 1 ) = - (y^2 -2y +1) - (-1) = -(y - 1)^2 + 1

OpenStudy (idealist10):

So if u=y-1, then du=dy, then what?

OpenStudy (anonymous):

Changing to \(u\), you get \[\int(u+1)\sqrt{1-u^2}\,du\] Try a trigo sub, \(u=\sin t\).

OpenStudy (idealist10):

Thank you guys~.